In a mass spectrometer chlorine ions of mass 35u and charge +5e are emitted from a source and accelerated through a potential difference of 250 kV. They then enter a region with a magnetic field which is perpendicular to their original direction of motion. The chlorine ions exit the spectrometer after being bent along a path with radius of curvature 3.5 m. What is the value of the magnetic field? (u = 1.66 ´ 10–27 kg, e = 1.6 ´ 10–19C)

Answer Choices:

A) 3.9 ´ 10–2 T

B) 2.5 ´ 10–2 T

C) 3.1 ´ 10–2 T

D) 5.4 ´ 10–2 T

E) none of the above
Answer #1

Speed of chlorine particles will be:

KE = dU

0.5*m*v^2 = q*V

v = sqrt (2*q*V/m)

v = sqrt (2*5*1.6*10^-19*250*10^3/(35*1.67*10^-27))

v = 2.616*10^6 m/sec

Now using force balance

Fc = Fm

m*v^2/r = qvB

B = m*v/qr

B = 35*1.67*10^-27*2.616*10^6/(5*1.6*10^-19*3.5)

B = 5.46*10^-2 T

Correct option is D.